OpenStudy (df001):

help with 7x^3y(3x^-1y^5/9x^3y^-2)^-3

OpenStudy (df001):

\[7x ^{3}y(\frac{ 3x ^{-1} y^5 }{ 9x ^{3}y ^{-2} })^{-3}\]

OpenStudy (df001):

I am stuck on this question and I need help before an hour for class

OpenStudy (mathstudent55):

Is this the problem? \(\large 7x^3y \left( \dfrac{3x^{-1}y^5}{9x^3y^-2} \right) ^{-3}\)

OpenStudy (df001):